Evaluate the limits. Applying L'Hopital's Rule Yield
lim┬(x→0)⁡〖1/x^3 〗 ∫_0^x▒〖t^2/(t^4+1) dt〗
1
Expert's answer
2020-09-09T17:25:51-0400

"\lim\limits_{x\rarr0}" "\frac{1}{x^3}" "\int_{0}^x" "\frac{t^2}{t^4+1}""dt" = "\lim\limits_{x\rarr0}" "\frac{1}{3x^2}""\frac{x^2}{x^4+1}" = "\lim\limits_{x\rarr0}" "\frac{1}{3(x^4+1)}" = "\frac{1}{3}"


The limit of the initial fraction equals the limit of the fraction with derivative of numerator and derivative of denominator. Derivative of denominator is

"\frac{d}{dx}" "(x^3)" = "3x^2" .

Derivative of numerator is

"\frac{d}{dx}" "\int_{0}^x" "\frac{t^2}{t^4+1}dt" = "\frac{d}{dx}" "(\Phi(x)-\Phi(0))"= "\frac{x^2}{x^4+1}" .

Here "\Phi(x)" is antiderivative of "\frac{t^2}{t^4+1}" in point "x".